I am totally having no luck with setting up and solving the following word problem, can somebody help, please?
1. An administrator of a web site is told that a new server can handle 11,000 "hits" per second. The web site is currently experiencing a peak demand of about 85,000 hits per second; but every month, the peak demand increases by 3500 hits per second. Use a proportion equation to determine how many new servers the administrator should buy to address expected traffic for the next 18 months.

Answer by stanbon(75887)

An administrator of a web site is told that a new server can handle 11,000 "hits" per second. The web site is currently experiencing a peak demand of about 85,000 hits per second; but every month, the peak demand increases by 3500 hits per second. Use a proportion equation to determine how many new servers the administrator should buy to address expected traffic for the next 18 months.
-----------------------------------------------
Increase in hits after 18 months = 3500*18 = 63000
-------------------------
Peak # of hits in 18 months: 85000 + 63000 = 148000
-------------------------
1/11000 = x/148000
x = 148000/11000 = 13.45
Rounding up the number of servers needed is 14
=====================================================
Cheers,
Stan H.